i need to find the perimeter of a rectangle here is the equation x+y=1350 and x/3=y

Answer by ewatrrr(24785)

 
Hi,

Assumming x and y are the lengths of the sides of this rectangle
Given: x+y=1350 and x/3=y
Perimeter = 2x + 2y = 2(x+y) = 2·1350 = 2700 units